b=4ft,h=9ft how do you solve this it's hard?
Answer by bucky(2097)

The formula for the area of a parallelogram is:
.
Area = b*h
.
All you have to do now is substitute into this equation the values of b and h that you are
given in the problem. The problem says that b = 4 ft and h = 9 ft. So the area of this
parallelogram is:
.
A = 4 * 9 = 36 sq ft
.
Hope that this helps you to understand how to get the area of any parallelogram if you are
given the measure of the base and the measure of the height.